Hi,
I was wondering if you could help me solve this system of equation algebraically.
Question: Solve the following systems of equations algebraically and check.
2/3x + y = 13
-x + 2y + 5
I am having truble understanding what to do since it has a fraction. I keep getting decimals and when I check the answers are not coming out correct.
Thanks for all your help. It is greatly appreciated.

Answer by mananth(16949)

2/3x + y = 13 , -x + 2y + 5
Whenever you get a fraction in an equation convert the fraction into a whole number by multiplying by the denominator or multiply by the LCD iof both are fractions.
2x/3+y = 13
multiply by 3
2x+3y= 39 ...............1
-x+2y= 5.................2 ( hope this is correct equation)
Eliminate y
multiply (1)by -2
Multiply (2) by 3
-4x-6y=-78
-3x+6y= 15
Add the two equations
-7x=-63
/-7
x=9
plug value of x in (1)
2x+3y=39
18+3y=39
3y=39-18
3y=21
y=7